# A Bias-Correction Decentralized Stochastic Gradient Algorithm with Momentum Acceleration

### v1

@@ 5 Conclusions
This paper focuses on the acceleration algorithm of decentralized stochastic gradient descent in sparse communication networks and data heterogeneity. By incorporating the idea of momentum acceleration into the bias-correction decentralized stochastic gradient algorithm, we propose an algorithm that can reach consensus faster when the data is heterogeneous. We prove theoretically that under non-convex and PL conditions, the proposed algorithm possesses at least comparable convergence property with the original algorithm, so it can eliminate the effect of data heterogeneity. However, concerning momentum SGD, the present analysis can only yield a convergence rate comparable to the original algorithm, and the specific principle of its acceleration is not yet clear, so how to construct the principle of distributed momentum acceleration, or more generally, centralized momentum SGD, remains to be further studied.
@@ Appendix E Simulations and Real-Data Analysis
In this section, we demonstrate the advantages of our algorithm through simulations. We compare our algorithm with the following two classes of algorithms: (i) momentum-free bias-correction algorithm (DSGT and ED/D2), and (ii) momentum-based algorithm (DmSGT-HB, DecentLaM, and Quasi-Global). The first type of algorithm is used to show that our algorithm accelerates the convergence of the original one, and the second type of algorithm is used to show that our algorithm has stronger adaptability to data heterogeneity and network sparsity. In the following three parts, we will simulate the square loss function, the general strongly convex objective function, and the non-convex objective function respectively.
For the double stochastic communication matrix $\mathbf{W}=(w_{ij})_{1\leq i,j\leq n}$. We model the sparsity of the communication network using a ring graph, where each agent is arranged in a circular formation and can only communicate with its two immediate neighbors. We define the weights as follows: $w_{i,i}=1/2,w_{i,i+1}=w_{i,i-1}=1/4$ and $w_{i,j}=0$ otherwise. In this context, we treat $-1$ as $n$ and $n+1$ as $1$ to account for the cyclic nature of the graph. We can check that the spectral gap of the above ring graph satisfies $1-\lambda=\mathcal{O}(1/n^{2})$. Therefore, with an increasing number of agents, the sparsity of the network will increase significantly.
@@ E.1 Quadratic Loss
We first consider a simple linear regression problem to visualize problems caused by network sparsity and data heterogeneity. Here, we consider the loss functions defined as
@@
where $i=1,\cdots,n$, $\mathbf{x}\in\mathbb{R}^{d},\mathbf{y}_{i}\in\mathbb{R}^{p}$, and $\mathbf{A}_{i}=(\mathbf{a}_{i1},\cdots,\mathbf{a}_{ip})^{\top}\in\mathbb{R}^{p\times d}$ represents the corresponding covariates on node $i$ and we choose $d=10$ and $p=20$. The covariates $\mathbf{a}_{i,j}$ are i.i.d. generated by $\mathcal{N}(0,\mathbf{I}_{d})$. And the local optimum parameter $\mathbf{x}_{i}^{\star}$ generated as follows. First, let $\mathbf{u}_{i}$ be i.i.d. drawn from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. Finally, let the exact minimizer $\mathbf{x}^{\star}=\left(\sum_{i}\mathbf{A}_{i}^{\top}\mathbf{A}_{i}\right)^{-1}\left(\sum_{i}\mathbf{A}_{i}^{\top}\mathbf{A}_{i}\mathbf{u}_{i}\right)$. Then let $\mathbf{x}_{i}^{\star}=\mathbf{x}^{\star}+(\mathbf{u}_{i}-\mathbf{x}^{\star})/c$. Thus we can control the data heterogeneity $\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\mathbf{x}^{\star})\|^{2}$ by $c$. Then $\mathbf{y}_{i}$ are generated as follows, $\mathbf{y}_{i}^{(t)}=\mathbf{A}_{i}\mathbf{x}_{i}^{\star}+\bm{\epsilon}_{i}^{(t)}$, where $\bm{\epsilon}_{i}^{(t)}$ are i.i.d. drawn from $\mathcal{N}(\mathbf{0}_{p},\sigma^{2}\mathbf{I}_{p})$.
With $n=32$ ($\lambda=0.99$) and $\sigma^{2}=0.05$, we experimented 20 times for various values of $\zeta^{2}$, and the results are presented in Figure 1.
@@
The experimental results indicate that all algorithms can converge linearly to a neighborhood of optimal values. However, as data heterogeneity increases, other momentum-based algorithms(DmSGD, Quasi-Global, and DecentLaM) become trapped in heterogeneous regions. Although the DSGT-HB algorithm effectively eliminates heterogeneity errors, its convergence rate is impacted by the original DSGT algorithm, resulting in momentum acceleration failure during the gradient consensus phase. In contrast, the EDM method retains the beneficial properties of the ED/D2, and the introduction of momentum improves the algorithm’s convergence rate.
@@ E.2 General Strongly Convex Loss
We consider the $\ell_{2}$-regularized logistic regression problem to investigate the algorithm’s convergence in a general strongly convex case. Let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ be the local parameter of node $i$, and $\bm{\xi}_{i}=\{(\mathbf{u}_{ij}^{\top},v_{ij})^{\top}\}_{j=1,\cdots,m}$, where $u_{ij}\in\mathbb{R}^{d}$ is the covariate and $v_{ij}\in\{-1,1\}$ is the response. Given $\mathbf{x}_{i}$ and $\mathbf{u}_{ij}$, the response $v_{ij}$ is generated by $v_{ij}=1$ with probability $1/(1+\exp(-\mathbf{x}_{i}^{\top}\mathbf{u}_{ij}))$ and $v_{ij}=-1$ otherwise. The loss function at agent $i$ can be expressed as
@@
It can be verified that the loss function is $\mu$-strongly convex.
@@
In our approach, we utilize the full batch gradient $\nabla F(\mathbf{x}_{i},\bm{\xi}_{i})$ at every iteration. However, since we are using full batch samples, this results in no randomness. Thus, we add an additional i.i.d. noise term $\mathbf{s}_{i}^{(t)}$ drawn from $\mathcal{N}(0,\sigma_{s}^{2}I_{d})$ to $\nabla F(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i})$. And use $\nabla F(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i})+\mathbf{s}_{i}^{(t)}$ as the stochastic gradient. This method is similar to the one used in Alghunaim and Yuan (2022) to control the influence of gradient noise.
@@
For our simulations, we set $d=20$ and $m=2000$, using the $\ell_{2}$-norm of the gradient $\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\|_{2}^{2}$ as a measure of convergence. The simulation is conducted 20 times on a ring graph with $n=32$ nodes. We compare our method against DSGT, and ED/D2, DSGT-HB and DmSGD. Figure 2 represents the result.
The simulation results demonstrate that our method can improve the performance of ED/D2, and achieve convergence to a narrower region of global optimal values, with minimal influence from data heterogeneity.
@@ E.3 Non-Convex Loss
Finally, we consider the non-convex condition. We deploy our method in a classification problem with deep neural networks and compare it with the aforementioned algorithms.
We use VGG-11 as the backbone network to train the classification task of the CIFAR10 dataset consisting of $N=50,000$ images of 10 labels, such as cats, airplanes, etc. The photos are $32\times 32$ size with three color channels. The criterion is the cross entropy. To quantitatively analyze the heterogeneity of the data in this dataset, we introduce the Dirichlet distribution, which is widely used to model heterogeneous scenarios for classification problems (Yurochkin et al., 2019; Lin et al., 2021). We generate $\mathbf{p}_{k}=(p_{k1},\cdots,p_{kn})$ for $k=1,\cdots,10$ from $\mathrm{Dirichlet}(\phi,\cdots,\phi)$, allocating $p_{ki}$ proportion of samples with label $k$ to agent $i$. The Dirichlet parameter $\phi$ is used to control the weights assigned to different categories of data across devices, serving as a measure of data heterogeneity. Specifically, a smaller value of $\phi$ indicates greater heterogeneity among the samples allocated to different agents. Our simulation for this part is primarily based on the code provided in Vogels et al. (2021).

### v2

@@ 5 Conclusions
This paper focuses on the acceleration algorithm of decentralized stochastic gradient descent in sparse communication networks and under conditions of data heterogeneity. By incorporating the idea of momentum acceleration into the bias-correction decentralized stochastic gradient algorithm, we propose a new algorithm that achieves consensus more rapidly in the presence of heterogeneous data. We theoretically demonstrate that, under non-convex and PL conditions, the proposed algorithm exhibits at least comparable convergence properties to the original algorithm, effectively mitigating the impact of data heterogeneity.
However, the specific mechanisms of momentum acceleration remain unclear. Therefore, further investigation is necessary to establish a robust framework for distributed momentum acceleration and, more broadly, for centralized momentum SGD.
@@ Appendix E Simulations and Real-Data Analysis
In this section, we demonstrate the advantages of our algorithm through simulations. We compare our algorithm with the following two classes of algorithms: (i) momentum-free bias-correction algorithm (DSGT and ED/D2), and (ii) momentum-based algorithm (DmSGT-HB, DecentLaM, and Quasi-Global). The first type of algorithm is used to show that our algorithm accelerates the convergence of the original one, and the second type of algorithm is used to show that our algorithm has stronger adaptability to data heterogeneity and network sparsity. In the following three parts, we will simulate the square loss function, the general strongly convex objective function, and the non-convex objective function, respectively.
For the double stochastic communication matrix $\mathbf{W}=(w_{ij})_{1\leq i,j\leq n}$, we model the sparsity of the communication network using a ring graph, where each agent is arranged in a circular formation and can only communicate with its two immediate neighbors. We define the weights as follows: $w_{i,i}=1/2,w_{i,i+1}=w_{i,i-1}=1/4$ and $w_{i,j}=0$ otherwise. In this context, we treat $-1$ as $n$ and $n+1$ as $1$ to account for the cyclic nature of the graph. It can be verified that the spectral gap of the above ring graph satisfies $1-\lambda=\mathcal{O}(1/n^{2})$. Consequently, as the number of agents increases, the sparsity of the network significantly increases.
@@ E.1 Quadratic Loss
We first consider a simple linear regression problem to visualize the issues caused by network sparsity and data heterogeneity. The loss functions are defined as follows:
@@
where $i=1,\cdots,n$, $\mathbf{x}\in\mathbb{R}^{d},\mathbf{y}_{i}\in\mathbb{R}^{p}$, and $\mathbf{A}_{i}=(\mathbf{a}_{i1},\cdots,\mathbf{a}_{ip})^{\top}\in\mathbb{R}^{p\times d}$ represents the corresponding covariates at node $i$, with the covariates $\mathbf{a}_{i,j}$ being i.i.d. samples generated from $\mathcal{N}(0,\mathbf{I}_{d})$. Here, we choose $d=10$ and $p=20$. The local optimum parameter $\mathbf{x}_{i}^{\star}$ generated as follows: first, let $\mathbf{u}_{i}$ be i.i.d. samples drawn from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. The exact minimizer is given by $\mathbf{x}^{\star}=\left(\sum_{i}\mathbf{A}_{i}^{\top}\mathbf{A}_{i}\right)^{-1}\left(\sum_{i}\mathbf{A}_{i}^{\top}\mathbf{A}_{i}\mathbf{u}_{i}\right)$. We then define $\mathbf{x}_{i}^{\star}=\mathbf{x}^{\star}+(\mathbf{u}_{i}-\mathbf{x}^{\star})/c$, allowing us to control the data heterogeneity $\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\mathbf{x}^{\star})\|^{2}$ through the parameter $c$. The responses $\mathbf{y}_{i}$ are generated as follows, $\mathbf{y}_{i}^{(t)}=\mathbf{A}_{i}\mathbf{x}_{i}^{\star}+\bm{\epsilon}_{i}^{(t)}$, where $\bm{\epsilon}_{i}^{(t)}$ are i.i.d. drawn from $\mathcal{N}(\mathbf{0}_{p},\sigma^{2}\mathbf{I}_{p})$.
With $n=32$ ($\lambda=0.99$) and $\sigma^{2}=0.05$, we conducted 20 experiments for various values of $\zeta^{2}$, and the results are presented in Figure 1.
@@
The experimental results indicate that all algorithms can converge linearly to a neighborhood of optimal values. However, as data heterogeneity increases, other momentum-based algorithms (DmSGD, Quasi-Global, and DecentLaM) become trapped in heterogeneous regions. Although the DSGT-HB algorithm effectively eliminates heterogeneity errors, its convergence rate is impacted by the original DSGT algorithm, resulting in momentum acceleration failure during the gradient consensus phase. In contrast, the EDM method retains the beneficial properties of the ED/D2, and the introduction of momentum improves the convergence rate.
@@ E.2 General Strongly Convex Loss
We consider the $\ell_{2}$-regularized logistic regression problem to investigate the algorithm’s convergence in a general strongly convex case. Let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ be the local parameter of node $i$, and $\bm{\xi}_{i}=\{(\mathbf{u}_{ij}^{\top},v_{ij})^{\top}\}_{j=1,\cdots,m}$, where $u_{ij}\in\mathbb{R}^{d}$ is the covariate and $v_{ij}\in\{-1,1\}$ is the response. Given $\mathbf{x}_{i}$ and $\mathbf{u}_{ij}$, the response $v_{ij}$ is generated such that $v_{ij}=1$ with probability $1/(1+\exp(-\mathbf{x}_{i}^{\top}\mathbf{u}_{ij}))$ and $v_{ij}=-1$ otherwise. The loss function at agent $i$ can be expressed as
@@
It can be verified that this loss function is $\mu$-strongly convex.
@@
We utilize the full batch gradient $\nabla F(\mathbf{x}_{i},\bm{\xi}_{i})$ at every iteration. However, since we are using full batch samples, this results in no randomness. Thus, we add an additional i.i.d. noise term $\mathbf{s}_{i}^{(t)}$ drawn from $\mathcal{N}(0,\sigma_{s}^{2}I_{d})$ to $\nabla F(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i})$. And use $\nabla F(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i})+\mathbf{s}_{i}^{(t)}$ as the stochastic gradient. This method is similar to the one used in Alghunaim and Yuan (2022) to control the gradient noise.
For our simulations, we set $d=20$ and $m=2000$, using the $\ell_{2}$-norm of the gradient $\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\|_{2}^{2}$ as a measure of convergence. The simulation is conducted 20 times on a ring graph with $n=32$ nodes. We compare our method against DSGT, ED/D2, DSGT-HB, and DmSGD. The results are presented in Figure 2.
@@
The simulation results demonstrate that our method enhances the performance of ED/D2, achieving convergence to a narrower region of global optimal values while experiencing minimal influence from data heterogeneity.
@@ E.3 Non-Convex Loss
Finally, we consider the non-convex condition. We use VGG-11 as the backbone network to train the classification task of the CIFAR10 dataset consisting of $N=50,000$ images of 10 labels, such as cats, airplanes, etc. The photos are $32\times 32$ size with three color channels. The criterion is the cross entropy. To quantitatively analyze the heterogeneity of the data in this dataset, we introduce the Dirichlet distribution, which is widely used to model heterogeneous scenarios for classification problems (Yurochkin et al., 2019; Lin et al., 2021). We generate $\mathbf{p}_{k}=(p_{k1},\cdots,p_{kn})$ for $k=1,\cdots,10$ from $\mathrm{Dirichlet}(\phi,\cdots,\phi)$, allocating $p_{ki}$ proportion of samples with label $k$ to agent $i$. The Dirichlet parameter $\phi$ is used to control the weights assigned to different categories of data across devices, serving as a measure of data heterogeneity. Specifically, a smaller value of $\phi$ indicates greater heterogeneity among the samples allocated to different agents. Our simulation for this part is primarily based on the code provided in Vogels et al. (2021).